patrick are 1/4 of the cake that was left.
there was 2/3 of the cake left. what part of the cake did patrick eat?

Answer by Mathtut(3670)

if there was 2/3 of the cake left and patrick ate 1/4 of that left over portion
:
then we have:
:
1/4 of 2/3 which is (1/4)(2/3)=2/12=1/6
:
therefore Patrick ate 1/6 of the cake